OpenStudy (anonymous):

Three horses, call them A, B, and C, are going to race against each other. The probability that A will win the race is 1⁄3 and the probability that B will win the race is 1⁄4 . What is the probability that C will win the race?

OpenStudy (anonymous):

Law to total probability: The sum of the probabilities of each outcome adds up to 1. So 1/3 + 1/4 + pobability C wins = 1

OpenStudy (anonymous):

P(C) = 1 - P(A) - P(B) = 5/12